Solving Exponential Equations Practice:

  1. 2x+ 3 = 11
  2. 2x+3 = 16
  3. 3x+1 - 7 = 20
1
Expert's answer
2021-05-18T13:38:13-0400

1.

"2^x +3 = 11 \\

2^x = 11 -3 \\

2^x = 8 \\

2^x = 2^3 \\

x = 3"

2.

"2^{x+3} = 16 \\

2^{x+3} = 2^4 \\

x+3 = 4 \\

x = 4-3 \\

x=1"

3.

"3^{x+1}-7 = 20 \\

3^{x+1} = 20 +7 \\

3^{x+1} = 27 \\

3^{x+1} = 3^3 \\

x+1 = 3 \\

x = 3-1 \\

x=2"